Jona is selling handmade specialty shirts at county fairs and farmer’s markets. The variable cost for each shirt is $13, and he sells them for $74. The cost to rent a booth at both the county fair and the farmer’s market is $205.


Q1: How many shirts must Jona sell to break even?
Q2: If Jona’s variable cost changes to $70 per unit, what would his break-even point be?
Answer #1

FORMULA:

BEQ = FIXED COST / PRICE PER UNIT - VARIABLE COST

FIXED COST = 205

VARIABLE COST = 13

PRICE PER UNIT = 74

BREAK EVEN UNITS = 205 / (74 - 13) = 3.36

Or 3 or 4 units depending on how we round off the calculation.


2. FIXED COST = 205

VARIABLE COST = 70

PRICE PER UNIT = 74

BREAK EVEN UNITS = 205 / (74 - 70) = 51.25

Or either 51 or 52 units.